# A Brain-Inspired Regularizer for Adversarial Robustness

### v1

@@ 2 Related works
Random noise attacks involve applying noise sampled from known distributions (e.g., Gaussian, Uniform, and Salt and Pepper) to an input - see Appendix 7.1. Common corruptions correspond to distortions that can be found in real life computer vision applications (eg : motion blur) [8]. A transfer-based attack involves finding adversarial perturbations for a substitute model (an unregularized model in our case) and applying them to a target model. Evaluating robustness on transfer-based attacks is crucial because adversarial examples crafted for one model can also mislead another distinct model [13]. We find these perturbations by applying the Fast Gradient Sign Method (FGSM) [14] to the substitute model. Such adversarial samples are computed as :
@@
where $x_{adv}$ is the adversarial example, $x$ denotes the original input image, $y$ denotes the original input label, $\theta$ denotes the model parameters and $L$ is the loss. A decision based attack is an attack which solely depends on the final model’s decision. Evaluating robustness on them is key as they are applicable to real world black box models. Precisely, we evaluate robustness on a Boundary Attack introduced by Brendel et al. [9]. This attack starts from a large adversarial perturbation, and seeks to reduce the perturbation while remaining adversarial.
Adversarial training as a defense to adversarial attacks. As adversarial attacks have advanced, corresponding defenses have also been developed to secure against them [14, 15, 16, 17]. A common defense strategy involves augmenting each batch of training data with adversarial examples, a technique known as adversarial training [14, 12]. A widely-used method for generating these adversarial examples is the Projected Gradient Descent (PGD) attack [12], which is a multi-step variant of the Fast Gradient Sign Method (FGSM) attack. PGD is popular due to its effectiveness in creating challenging adversarial examples, thereby enhancing the model’s resilience against attacks.
Neural regularization. Recent work showed that jointly training a deep network to perform image classification while steering it towards having brain-like representations can improve the model’s robustness to adversarial attacks [1, 5]. This is achieved by introducing a penalty term in the loss function acting directly on image representations [1] or activations [5] at different network depths. Such a process is referred to as neural regularization. For instance, [1] used a neural regularizer to drive a CNN to aligh its representational similarities [7] with those of mouse primary visual cortex (V1). Later on, Safarani et al. [5] used a neural regularizer to drive a CNN towards predicting neural activity in macaque primary visual cortex (V1) in response to the same natural stimuli. The key bottleneck of such defenses is their reliance on the measurement of large scale neural recordings.
@@ 3 A neural representational similarity regularizer
and $\alpha$ is a parameter that sets the overall regularization strength. $S_{ij}^{\text{target}}$ in eq. (3) is the target’s pairwise cosine similarity between the representations of images $i$ and $j$. $S^{\text{CNN}}_{ij}$ measures the similarity between the representations of images $i$ and $j$ in a CNN. We compute it following the approach of Li et al. [1]. We combine feature similarities from a selection of $K$ equally spaced convolutional layers and average the results through a trainable weights $\gamma_{l}$, where $l$ is the layer number. The latter are the output of a softmax function meaning $\sum_{l}\gamma_{l}=1$ and $\gamma_{l}\geq 0$. Overall,
@@ 4 A pixel similarity based regularizer
We inspect the similarity loss $L_{\text{sim}}$ (eq.(3)) introduced in [1] to extract its underlying strength. Our goal is to formulate a method that can bypass the use of neural recordings which can be costly.
Since the primary visual cortex (V1), where the neural recordings come from, is the first visual processing area in the cortex, we inspect the correlation between the neural representational similarity and image pixel similarities (computed as described in Appendix 7.1). We observe that there is a high correlation between the two as shown in Fig. 2, left panel. Thus, we investigate the effect of using pixel similarities as target similarities in $L_{\text{sim}}$ instead of similarities obtained from neural recordings. To compare both approaches, we replicate the experimental setup in Li et al. [1]. We train a ResNet to classify grayscale CIFAR-10 and CIFAR-100, and use grayscale images from ImageNet data, as the regularization dataset (same datasets used in [1]). However, we differ from [1] in our choice of $S_{ij}^{\text{target}}$: we set $S_{ij}^{\text{target}}$ in eq. (3) to $S_{ij}^{\text{pixel}}$, where $S_{ij}^{\text{pixel}}$ is computed as the pixel cosine similarity between images which are flattened, mean subtracted and normalized. After training, we observe that the regularized model exhibits some enhancement in robustness, however this enhancement is not consistent across different image perturbations and adversarial attacks. For example, we see a modest enhancement in robustness to Gaussian noise (Fig. 1(a)), and decision-based Boundary Attack (Fig. 1(c)), but this was not the case for transferred FGSM attack (Fig. 1(b)). For Uniform noise and Salt and Pepper noise, we see an enhancement in robustness at large noise levels (see Fig. 12 in 7.2). In Fig. 1 we also show the performance of the model of Li et al. [1], which is regularized using neural representational similarities, for direct comparison.
Figure 1: ResNet18 classifying grayscale CIFAR-10 is regularized using ImageNet ($S^{\text{pixel}}$) or a neural predictive model trained on ImageNet as in [1]. (a) Robustness to Gaussian Noise, (b) Transferred FGSM [14] perturbations and (c) to a decision-based Boundary Attack [9] as shown by the difference in retrieved perturbation sizes retrieved. Neural and pixel regularized models use $\alpha=10$. Error shades/bars represent the SEM across 7 seeds per model - except for neural models where we had access to 5 models. (Details on the experimental setup can be found in Appendix 7.1)
@@
If we go back and visually examine the last two panels in Fig.2, we observe that the image pixel and neural representational similarity matrices have similar patterns, however the pattern is more enhanced in the neural representational similarity matrix. Thus, at the first stage of cortical visual processing, the brain seems to roughly preserve an underlying structure of the image pixel similarities but amplify it. Based on this observation, we define a new target similarity $S^{Th}$, such that
@@
where, $Th\in(0,1)$ is a tunable thresholding hyperparameter. That is, we set $S_{ij}^{\text{target}}$ to be $S_{ij}^{Th}$. We note that in practice we don’t use exactly $1$ and $-1$ in eq. (5) by a very small number $\epsilon$ because of the arctanh function in eq.(3). Finally, we note that even though we use $L_{\text{sim}}$ as in (3), the application of the arctanh function is not necessary in this case. The intuition behind this regularization term is as follow: it is constraining the lower layers in the network (as we show in Appendix 7.6) to have identical representations for image pairs that are close in pixel space, measured by the cosine similarity; hence viewing those images pairs as adversarial versions of each other. At the same time, it is pushing images farther away in pixel space to have orthogonal representations. In 7.7, we show the contribution of each term of $S^{Th}$ on robustness. In section 5.2 we propose how to select the hyperparameter $Th$ and the regularization hyperparameter $\alpha$.
Key advantages for this regularization method are 1) it does not require access to large scale neural recordings, 2) it relies on the original datasets, and does not require the introduction of data distortions or augmentations during training, 3) it is computationally inexpensive, as we show later in Section 5.6 .
@@ 5.1 Robustness to adversarial attacks
We first test the robustness of regularized models using grayscale CIFAR-10 and grayscale ImageNet as classification and regularization datasets respectively, as used in Li et al. [1]. We first show robustness to Gaussian noise perturbations. We find that regularized models exhibit a substantial increase in robustness when compared to unregularized models as seen in Fig.3, left panel. They also show a similar performance to neural regularized models [1] (Fig.3, left panel). The robustness of models regularized using $S^{Th}$, to Uniform and Salt and Pepper perturbations can be found in appendix 7.3.
We then test robustness to stronger black box attacks, particularly, to transferred FGSM [14] perturbations from an unregularized model, and decision-based Boundary Attack [9] (Section 2). We observe an increase in robustness to both attacks (Fig.3, center and right panels). Note that for decision-based Boundary attack, the larger the perturbation size between the adversarial input and the original image, the better in terms of robustness. Again, we observe that models regularized using $S^{th}$ perform similar to those regularized using neural data [1] (Fig. 3, center and right panels).
@@
Figure 3: Robustness to Gaussian noise (left), transferred FGSM [14] (center), and decision-based Boundary Attack [9] (right). A ResNet18 is trained to classify grayscale CIFAR-10 and is regularized on grayscale images from ImageNet dataset. Results for different regularization targets are shown : $S^{pixel},S^{Th}$ and neural based targets as in [1]. For the decision-based Boundary Attack, we compute the median squared $L_{2}$ perturbation size per pixel, averaged across 1000 images, and 5 repeats. Error shades represent the SEM across seven seeds per model.
@@ 5.2 Hyperparameter selection and consistent behavior across attacks
We can visualize the performance of a model by plotting $\frac{R_{0}}{U_{0}}$ vs $\frac{R_{D}}{U_{D}}$) for each $\alpha$, $Th$ pair. This allows the user to select the hyperparameter pair based on the selection criterion that they choose. In Fig.4 we show the above plot for different adversarial attacks (the gray shaded planes).
@@
Figure 4: Behavior across multiple black box attacks and hyperparameters $(\alpha,Th)$ choices. Models are trained to classify grayscale CIFAR-10 and regularized on grayscale images from ImageNet dataset. Different planes show results for different black box attacks. In each plane, the region shaded in blue represents the region of ’acceptable’ models, which we’ve taken here to be $R_{D}/U_{D}\geq 1$ and $R_{0}/U_{0}\geq 0.9$ – a criteria that can be adjusted as needed. $U_{D},R_{D}$ are computed at $\epsilon=0.1$ for random attacks and $\epsilon=0.02$ for the transferred FGSM [14] attack. Mean metrics across 7 seeds per model are displayed.
@@ 5.3 Robustness across datasets combinations
Figure 5: Robustness of a ResNet18 trained to classify grayscale CIFAR-10 regularized on grayscale images from different datasets : grayscale CIFAR-10 (blue), CIFAR-100 (purple) or ImageNet (red). For the decision-based Boundary Attack, we compute the median $L_{2}$ perturbation size, averaged across 1000 images, and 5 repeats. Error shades/bars represent the SEM across seven seeds per model. The same ($\alpha$,$Th$) values were used in training all models i.e for all regularization datasets (see Appendix 7.5).
@@ 5.4 Robustness to common corruptions
Regularized models are also more robust than their unregularized counterpart on common corruptions. We evaluate regularized models on grayscale CIFAR-10-C dataset [8] which consists of grayscaled CIFAR-10 images with common corruptions that can be found on everyday computer vision application. Evaluating on common corruptions at different severity levels is critical as they simulate real world conditions. Fig. 6 shows the performance of a ResNet18 trained to classify CIFAR-10 regularized with grayscale images from ImageNet dataset vs unregularized model. Fig. 6 (left) shows the performance averaged over all 15 common corruptions at different severity levels, Fig. 6 (right) shows the robustness of unregularized and regularized models for the 15 individual corruptions present in CIFAR-10-C [8], at severity level 4.
Figure 6: Robustness to grayscale CIFAR-10-C Common Corruptions [8]. (left) We compute the regularized model accuracy on grayscale CIFAR-10-C for different severity levels, averaging across all 15 common corruptions present in CIFAR-10-C. (right) We show the robustness of unregularized and regularized models on 15 individual corruptions at severity 4. Error bars correspond to the SEM across seven seeds per model. Results are for a ResNet18 trained to classify grayscale CIFAR-10 regularized wih grayscale images from ImageNet dataset.
@@ 5.5 Frequency decomposition of adversarial perturbations and common corruptions
To understand the strengths and weaknesses of our regularization method, we investigate the frequency components present in the minimal perturbation required to flip the decision of unregularized and regularized models which we compute via a decision-based Boundary Attack [9]. We observe that models regularized using pixel-based similarities ($S^{Th}$) rely more on low frequency information than their unregularized counterparts (Fig. 7 center and right panels). We further evaluate our regularized model performance on grayscale CIFAR-10-C [8] following the approach described in [6], where we categorize the 15 corruptions in CIFAR-10-C into Low, Medium, and High frequency based on their spectra (see Fig. 7 left panel and Appendix Fig. 25). Results are shown for ResNet18 trained to classify CIFAR-10 and regularized using images from CIFAR-10, CIFAR-100 or imageNet datasets. Our results show that regularized models outperform unregularized ones, especially on high-frequency corruptions, confirming our findings. Such a reliance on low-frequency information has also been observed in models subjected to neural regularization as explained in [6].
Figure 7: Frequency perspective on robustness. The results are for ResNet18 trained to classify grayscale CIFAR-10 and regularized on grayscale images from different datasets: CIFAR-10, CIFAR-100 or ImageNet. (left) Robustness of regularized ResNet18 models evaluated on grayscale CIFAR-10-C at severity 4, categorized by the frequency range of each corruption. (center) Fourier power spectrum for the mean minimal corruption required to flip a model’s decision. (right) Radial Spectrum of minimal pertubation required to mislead models, as provided by a decision-based Boundary Attack [9] - using 10k steps. The error bars (left panel) and shaded areas (right panel) represent the SEM across seven and four seeds per model respectively.

### v2

@@ 2 Related works
Random noise attacks involve applying noise sampled from known distributions (e.g., Gaussian, Uniform, and Salt and Pepper) to an input, see Appendix 7.1. Common corruptions correspond to distortions that can be found in real life computer vision applications (eg : motion blur) [8]. A transfer-based attack involves finding adversarial perturbations for a substitute model (an unregularized model in our case) and applying them to a target model. Evaluating robustness on transfer-based attacks is crucial because adversarial examples crafted for one model can also mislead another distinct model [13]. We find these perturbations by applying the Fast Gradient Sign Method (FGSM) [14] to the substitute model. Such adversarial samples are computed as :
@@
where $x_{adv}$ is the adversarial example, $x$ denotes the original input image, $y$ denotes the original input label, $\theta$ denotes the model parameters and $L$ is the loss. A decision based attack is an attack which solely depends on the final model’s decision. Evaluating robustness on them is key as they are applicable to real world black box models. Precisely, we evaluate robustness on a Boundary Attack introduced by Brendel et al. [9]. This attack starts from a large adversarial perturbation, and seeks to reduce the perturbation while remaining adversarial. Note that for such attack, the larger the perturbation size between the final adversarial image and the original image, the better in terms of robustness.
Adversarial training as a defense to adversarial attacks. As adversarial attacks have advanced, corresponding defenses have also been developed to secure against them [14, 15, 16, 17]. A common defense strategy involves augmenting each batch of training data with adversarial examples, a technique known as adversarial training [14, 12]. A widely used method for generating these adversarial examples is the Projected Gradient Descent (PGD) attack [12], which is a multi-step variant of the Fast Gradient Sign Method (FGSM) attack. PGD is popular due to its effectiveness in creating challenging adversarial examples, thereby enhancing the model’s resilience against attacks.
Neural regularization. Recent work showed that jointly training a deep network to perform image classification while steering it towards having brain-like representations can improve the model’s robustness to adversarial attacks [1, 5]. This is achieved by introducing a penalty term in the loss function acting directly on image representations [1] or activations [5] at different network depths. Such a process is referred to as neural regularization. For instance, Li et al. [1] used a neural regularizer to drive a CNN to aligh its representational similarities [7] with those of mouse primary visual cortex (V1). Later on, Safarani et al. [5] used a neural regularizer to drive a CNN towards predicting neural activity in macaque primary visual cortex (V1) in response to the same natural stimuli. The key bottleneck of such defenses is their reliance on the measurement of large scale neural recordings.
@@ 3 A neural representational similarity regularizer
and $\alpha$ is a parameter that sets the overall regularization strength. $S_{ij}^{\text{target}}$ in eq. (3) is the target’s pairwise cosine similarity between the representations of images $i$ and $j$. $S^{\text{CNN}}_{ij}$ measures the similarity between the representations of images $i$ and $j$ in a CNN. We compute it following the approach of Li et al. [1]. We combine feature similarities from a selection of $K$ equally spaced convolutional layers and average the results through trainable weights $\gamma_{l}$, where $l$ is the layer number. The latter are the output of a softmax function meaning $\sum_{l}\gamma_{l}=1$ and $\gamma_{l}\geq 0$. Overall,
@@ 4 A pixel similarity based regularizer
We inspect the similarity loss $L_{\text{sim}}$ (eq. (3)) introduced in Li et al. [1] to extract its underlying strength. Our goal is to formulate a method that can bypass the use of neural recordings which can be costly.
Since the primary visual cortex (V1), where the neural recordings come from, is the first visual processing area in the cortex, we inspect the correlation between the neural representational similarity and image pixel similarities (computed as described in Appendix 7.1).
@@
We observe that there is a high correlation between the two as shown in Fig. 1, left panel. Thus, we investigate the effect of using pixel similarities as target similarities in $L_{\text{sim}}$ instead of similarities obtained from neural recordings. To compare both approaches, we replicate the experimental setup in Li et al. [1]. We train a ResNet18 to classify grayscale CIFAR-10, and use grayscale images from ImageNet data, as the regularization dataset (same datasets used in Li et al. [1]). However, we differ from Li et al. [1] in our choice of $S_{ij}^{\text{target}}$: we set $S_{ij}^{\text{target}}$ in eq. (3) to $S_{ij}^{\text{pixel}}$, where $S_{ij}^{\text{pixel}}$ is computed as the pixel cosine similarity between images which are flattened, mean subtracted and normalized. After training, we observe that the regularized model exhibits some enhancement in robustness, however this enhancement is not consistent across different image perturbations and adversarial attacks - as shown in Fig. 2.
Figure 2: ResNet18 classifying grayscale CIFAR-10 is regularized using ImageNet ($S^{\text{pixel}}$) or a neural predictive model trained on ImageNet as in [1]. (a) Robustness to Gaussian Noise, (b) Transferred FGSM [14] perturbations and (c) to a decision-based Boundary Attack [9] as shown by the difference in retrieved perturbation sizes retrieved. Neural and pixel regularized models use $\alpha=10$. Error shades/bars represent the SEM across 7 seeds per model - except for neural models where we had access to 5 models. (Details on the experimental setup can be found in Appendix 7.1).
For example, we see a modest enhancement in robustness to Gaussian noise (Fig. 2(a)), and decision-based Boundary Attack (Fig. 2(c)), but this was not the case for transferred FGSM attack (Fig. 2(b)). For Uniform noise and Salt and Pepper noise, we see an enhancement in robustness at large noise levels (see Fig. 12 in Appendix 7.2). In Fig. 2 we also show the performance of the model of Li et al. [1], which is regularized using neural representational similarities, for direct comparison.
If we go back and visually examine the last two panels in Fig. 1, we observe that the image pixel and neural representational similarity matrices have similar patterns, however the patterns are more enhanced in the neural representational similarity matrix. Thus, at the first stage of cortical visual processing, the brain seems to roughly preserve an underlying structure of the image pixel similarities but amplify it. Based on this observation, we define a new target similarity $S^{Th}$, such that
@@
where, $Th\in(0,1)$ is a tunable thresholding hyperparameter. That is, we set $S_{ij}^{\text{target}}$ to be $S_{ij}^{Th}$. We note that in practice we don’t use exactly $1$ and $-1$ in eq. (5) by a very small number $\epsilon$ because of the arctanh function in eq. (3). Finally, we note that even though we use $L_{\text{sim}}$ as in eq. (3), the application of the arctanh function is not necessary in this case. The intuition behind this regularization term is as follows: it is constraining the lower layers in the network (as we show in Appendix 7.6) to have identical representations for image pairs that are close in pixel space, measured by the cosine similarity; hence viewing those images pairs as adversarial versions of each other. At the same time, it is pushing images farther away in pixel space to have orthogonal representations. In Appendix 7.7, we show the contribution of each term of $S^{Th}$ on robustness. In Section 5.2 we propose how to select the hyperparameter $Th$ and the regularization hyperparameter $\alpha$.
Key advantages for this regularization method are 1) it does not require large scale neural recordings, 2) it relies on the original datasets, and does not require the introduction of data distortions or augmentations during training, and 3) it is computationally inexpensive, as we show in Section 5.6 .
@@ 5.1 Robustness to adversarial attacks
We test the robustness of regularized models using grayscale CIFAR-10 and grayscale ImageNet as classification and regularization datasets respectively, as used in Li et al. [1]. For Gaussian noise perturbations, we find that regularized models exhibit a substantial increase in robustness when compared to unregularized models as seen in Fig. 3, left panel. They also show a similar performance to neural regularized models [1] (Fig. 3, left panel). The robustness of models regularized using $S^{Th}$, to Uniform and Salt and Pepper perturbations are shown in Appendix 7.3.
Figure 3: Robustness to Gaussian noise (left), transferred FGSM [14] (center), and decision-based Boundary Attack [9] (right). A ResNet18 is trained to classify grayscale CIFAR-10 and is regularized on grayscale images from ImageNet dataset. Results for different regularization targets are shown : $S^{pixel},S^{Th}$ and neural based targets as in Li et al. [1]. For the decision-based Boundary Attack, we compute the median squared $L_{2}$ perturbation size per pixel, averaged across 1000 images, and 5 repeats. Error shades represent the SEM across seven seeds per model.
We also test robustness to stronger black box attacks, particularly, to transferred FGSM [14] perturbations from an unregularized model, and decision-based Boundary Attack [9] (Section 2). We observe an increase in robustness to both attacks (Fig. 3, center and right panels). Note that for decision-based Boundary attack, the larger the perturbation size between the adversarial input and the original image, the better in terms of robustness. Again, we observe that models regularized using $S^{Th}$ perform similar to those regularized using neural data [1] (Fig. 3, center and right panels).
@@ 5.2 Hyperparameter selection and consistent behavior across attacks
We can visualize the performance of a model by plotting $\frac{R_{0}}{U_{0}}$ vs $\frac{R_{D}}{U_{D}}$ for each $\alpha$, $Th$ pair. This allows the user to select the hyperparameter pair based on the selection criterion that they choose. In Fig. 4 we show the above plot for different adversarial attacks (the gray shaded planes).
Figure 4: Behavior across multiple black box attacks and hyperparameters $(\alpha,Th)$ choices. Models are trained to classify grayscale CIFAR-10 and regularized on grayscale images from ImageNet dataset. Different planes show results for different black box attacks. In each plane, the region shaded in blue represents the region of ’acceptable’ models, which we’ve taken here to be $R_{D}/U_{D}>1$ and $R_{0}/U_{0}\geq 0.9$ – a criteria that can be adjusted as needed. $U_{D},R_{D}$ are computed at $\epsilon=0.1$ for random attacks and $\epsilon=0.02$ for the transferred FGSM [14] attack. Mean metrics across 7 seeds per model are displayed. The ($\alpha,Th$) values that we use in our experiments are reported in Appendix 7.5.
@@ 5.3 Robustness across datasets combinations
Figure 5: Robustness of a ResNet18 trained to classify grayscale CIFAR-10 regularized on grayscale images from different datasets : grayscale CIFAR-10 (blue), CIFAR-100 (purple) or ImageNet (red). For the decision-based Boundary Attack, we compute the median $L_{2}$ perturbation size, averaged across 1000 images, and 5 repeats. Error shades/bars represent the SEM across seven seeds per model. The same ($\alpha$,$Th$) values are used in training all models i.e for all regularization datasets (see Appendix 7.5).
@@ 5.4 Robustness to common corruptions
Regularized models are also more robust than their unregularized counterpart on common corruptions. We evaluate regularized models on grayscale CIFAR-10-C dataset [8] which consists of grayscaled CIFAR-10 images with common corruptions that can be found on everyday computer vision application. Evaluating on common corruptions at different severity levels is critical as they simulate real world conditions. Fig. 6 shows the performance of a ResNet18 trained to classify grayscale CIFAR-10 regularized with grayscale images from ImageNet dataset vs unregularized model. Fig. 6 (left) shows the performance averaged over all 15 common corruptions at different severity levels, Fig. 6 (right) shows the robustness of unregularized and regularized models for the 15 individual corruptions present in CIFAR-10-C [8], at severity level 4.
Figure 6: Robustness to grayscale CIFAR-10-C Common Corruptions [8]. (left) We compute the regularized model’s accuracy on grayscale CIFAR-10-C for different severity levels, averaging across all 15 common corruptions present in CIFAR-10-C. (right) We show the robustness of unregularized and regularized models on 15 individual corruptions at severity 4. Error bars correspond to the SEM across seven seeds per model. Results are for a ResNet18 trained to classify grayscale CIFAR-10 regularized wih grayscale images from ImageNet dataset.
@@ 5.5 Frequency decomposition of adversarial perturbations and common corruptions
To understand the strengths and weaknesses of our regularization method, we investigate the frequency components present in the minimal perturbation required to flip the decision of unregularized and regularized models which we compute via a decision-based Boundary Attack [9]. We observe that models regularized using pixel-based similarities ($S^{Th}$) rely more on low frequency information than their unregularized counterparts (Fig. 7 center and right panels). We further evaluate our regularized model performance on grayscale CIFAR-10-C [8] following the approach described in [6], where we categorize the 15 corruptions in CIFAR-10-C into Low, Medium, and High frequency based on their spectra (see Fig. 7 left panel and Appendix Fig. 25). Results are shown for ResNet18 trained to classify CIFAR-10 and regularized using images from CIFAR-10, CIFAR-100 or ImageNet datasets. Our results show that regularized models outperform unregularized ones, especially on high-frequency corruptions, confirming our findings. Such a reliance on low-frequency information has also been observed in models subjected to neural regularization as explained in Li et al. [6].
Figure 7: Frequency perspective on robustness. The results are for ResNet18 trained to classify grayscale CIFAR-10 and regularized on grayscale images from different datasets: CIFAR-10, CIFAR-100 or ImageNet (left). Robustness of regularized ResNet18 models evaluated on grayscale CIFAR-10-C at severity 4, categorized by the frequency range of each corruption. (center) Fourier power spectrum for the mean minimal corruption required to flip a model’s decision. (right) Radial Spectrum of minimal pertubation required to mislead models, as provided by a decision-based Boundary Attack [9] - using 10k steps. The error bars (left panel) and shaded areas (right panel) represent the SEM across seven and four seeds per model respectively.